OpenStudy (anonymous):

What is the factored form of the expression? 4x^2-8ly^2

jimthompson5910 (jim_thompson5910):

Hint: \(\LARGE 4x^2 = (2x)^2\) and \(\LARGE 81y^2 = (9y)^2\) -------------------- So \(\LARGE 4x^2-81y^2\) is the same as \(\LARGE (2x)^2 - (9y)^2\)

jimthompson5910 (jim_thompson5910):

Tell me what you get or if you need more help.

OpenStudy (anonymous):

I really dont understand it at all

jimthompson5910 (jim_thompson5910):

Do you remember the difference of squares rule?

OpenStudy (anonymous):

ohh i get it now sorry :)

jimthompson5910 (jim_thompson5910):

that's ok, so what's the final answer?

OpenStudy (anonymous):

(2x-9y)^2

jimthompson5910 (jim_thompson5910):

close, but not quite

jimthompson5910 (jim_thompson5910):

the difference of squares rule is \(\LARGE a^2 - b^2 = (a-b)(a+b)\)

OpenStudy (anonymous):

what about (2x+9y)^2

jimthompson5910 (jim_thompson5910):

not that either

jimthompson5910 (jim_thompson5910):

compare \(\LARGE (2x)^2 - (9y)^2\) with \(\LARGE a^2 - b^2 = (a-b)(a+b)\)

jimthompson5910 (jim_thompson5910):

notice how a = 2x and b = 9y

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so \(\LARGE a^2 - b^2 = (a-b)(a+b)\) \(\LARGE (2x)^2 - (9y)^2 = (2x-9y)(2x+9y)\)

jimthompson5910 (jim_thompson5910):

since \(\LARGE 4x^2-81y^2 = (2x)^2 - (9y)^2\) This means \(\LARGE 4x^2-81y^2 = (2x-9y)(2x+9y)\)

OpenStudy (anonymous):

so its (2x-9y)(2x+9y)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thank you! :)

jimthompson5910 (jim_thompson5910):

sure thing